OpenStudy (anonymous):

using elimination solve the system of equations x+2y=10 x-y=4 i need to show all work please help

OpenStudy (cwrw238):

x + 2y = 10 x - y = 4 subtract: 3y = 6 y = 2 now substitute for y in second equation: x - 2 = 4 x = 4 + 2 = 6 solution set is {6, 2}

OpenStudy (anonymous):

thank you so much :)

OpenStudy (cwrw238):

yw